# Interval and $\ell$-interval rational parking functions

### v4

@@ Abstract.
1991 Mathematics Subject Classification:
Primary: 05A05; Secondary 05A15, 05A18, 05A19
1. Introduction
Throughout, we let $\mathbb{N}\coloneqq\{0,1,2,3,\ldots\}$, +$\mathbb{N}^{+}\coloneqq\{1,2,3,\ldots\}$, and for $n\in\mathbb{N}^{+}$, we let $[n]\coloneqq\{1,2,\ldots,n\}$. We begin by recalling a definition of parking functions, as introduced by Konheim and Weiss [24]. With $n\in\mathbb{N}^{+}$, a tuple $\alpha=(a_{1},a_{2},\ldots,a_{n})\in[n]^{n}$ is a parking function of length $n$ if its nondecreasing rearrangement $\alpha^{\prime}=(a_{1}^{\prime},a_{2}^{\prime},\ldots,a_{n}^{\prime})$ satisfies $a_{i}^{\prime}\leq i$ for all $i\in[n]$. For example, $\alpha=(3,1,1,3,4,5)$ is a parking function of length six as its nondecreasing rearrangement $\alpha^{\prime}=(1,1,3,3,4,5)$ satisfies the required inequality conditions. However, $\alpha=(2,3,3,3)$ is not a parking function, as its nondecreasing rearrangement $\alpha^{\prime}=(2,3,3,3)$ does not satisfy the inequality conditions: In particular, $a_{1}=2\nleq 1$. Let $\mathrm{PF}_{n}\subseteq[n]^{n}$ denote the set of parking functions of length $n$. Several authors have provided proofs that $|\mathrm{PF}_{n}|=(n+1)^{n-1}$, see [24, 28, 30].
@@ 1. Introduction
Figure 1. The parking outcome of the parking function $\alpha=(3,1,1,3,4,5)\in\mathrm{PF}_{6}$.
Parking functions have numerous applications and connections to many fields in mathematics, including computing volumes of flow polytopes, hyperplane arrangements, and the sorting algorithms [5, 6, 23]. Work on parking functions includes the study of discrete statistics [1, 18, 29], families of parking functions with particular desirable properties [2, 16], and many generalizations of parking functions [12, 13, 17, 22]. Our work is motivated by two generalizations of parking functions, which we refer to as “rational” parking functions and interval parking functions. We describe those next.
First we consider the situation in which there are $n\in\mathbb{N}^{+}$ cars and $m\geq n$ parking spots on a one-way street. In this case, a preference $\alpha$ is an element of $[m]^{n}$, and the set $\mathrm{PF}_{n,m}$ consists of the preferences allowing all $n$ cars to park under the standard parking procedure. This generalization of parking functions is known as $(m,n)$-parking functions. In the case of $n$ and $m$ being coprime, $(m,n)$-parking functions are referred to as rational parking functions, see [4]. In the current work, we simply refer to $(m,n)$-parking functions as rational parking functions without requiring $m$ and $n$ to be coprime. These parking functions were enumerated during the onset of parking function studies [24, Lemma 2] and, implicitly, in the statistical work of Pyke [28], giving
$|\mathrm{PF}_{n,m}|=\left(m+1-n\right)(m+1)^{n-1}.$ (1.1)
Another generalization of parking functions are interval parking functions which describe the case where cars have an interval of the street in which they want to park. Formally, interval parking functions are defined by a pair $(\alpha,\beta)$, where $\alpha=(a_{1},a_{2},\ldots,a_{n})\in\mathrm{PF}_{n}$ and $\beta=(b_{1},b_{2},\ldots,b_{n})\in[n]^{n}$, satisfy, for all $i\in[n]$, the inequality $a_{i}\leq b_{i}\leq n$. For $i\in[n]$, we call $a_{i}$ the initial preference of car $i$, $b_{i}$ the tolerance of car $i$ (as it is the highest numbered parking spot in which car $i$ would tolerate parking), and the set $\{a_{i},a_{i}+1,\ldots,b_{i}\}$ the preference interval for car $i$. Throughout, we refer to $\beta$ as the tolerance vector. If, for each $i\in[n]$, car $i$ is able to park in its preference interval following the standard parking procedure, then the pair $(\alpha,\beta)$ is called an interval parking function. We let $\mathrm{\mathrm{IPF}}_{n}$ denote the set of interval parking functions of length $n$. Colaric, DeMuse, Martin, and Yin established [12, Proposition 3.2]:
$|\mathrm{\mathrm{IPF}}_{n}|=n!(n+1)^{n-1}.$ (1.2)
@@
In the present work, we begin by generalizing interval parking functions to their rational analogue, which we call interval rational parking functions. Here, there are $n$ cars and $m\geq n$ parking spots on the street. In this case, we consider the pair of tuples $(\alpha,\beta)$, where $\alpha=(a_{1},a_{2},\ldots,a_{n})\in\mathrm{PF}_{n,m}$ and $\beta=(b_{1},b_{2},\ldots,b_{n})\in[m]^{n}$ satisfying $a_{i}\leq b_{i}\leq m$. If, for each $i\in[n]$, car $i$ is able to park in one of the spots numbered $a_{i},a_{i}+1,\ldots,b_{i}$ (following the standard parking procedure), then $(\alpha,\beta)$ is an interval rational parking function. We let $\mathrm{\mathrm{IPF}}_{n,m}$ denote the set of interval rational parking functions with $n$ cars and $m$ spots. For example, $((3,5,7,1,1),(3,6,8,1,4))\in\mathrm{\mathrm{IPF}}_{5,8}$ is a rational interval parking function with five cars and eight parking spots whose parking outcome is illustrated in Figure 2. However, $((3,5,7,1,1),(3,6,8,1,1))$ is not an interval rational parking function for any $\mathrm{\mathrm{IPF}}_{5,m}$ with $m\geq 8$, as car $5$ would fail to park in the first spot, which is the only spot in its preference interval. In our first result, we establish a formula for $|\mathrm{\mathrm{IPF}}_{n,m}|$ as a sum over rational parking functions involving the parking outcome, i.e. the order in which the cars park on the street (Theorem 3.4). Specializing $m=n$ recovers the formula in Equation (1.2) for the number of interval parking functions.
@@
Figure 2. The parking outcome of the interval rational parking function $((3,5,7,1,1),(3,6,8,1,4))\in\mathrm{\mathrm{IPF}}_{5,8}$.
@@
Fubini rankings and unit interval parking functions have received much recent attention in the literature. In [7, Theorem 2.9], the authors give a complete characterization of unit interval parking functions. In [10], the authors establish that unit interval parking functions with $k$ unlucky cars (cars that do not park in their preference) are in bijection with the $k$-dimensional faces of the permutohedron of order $n$. In [15], the authors consider the set of unit Fubini rankings, which is the intersection of the set of unit interval parking functions of length $n$ and the set of Fubini rankings. Through the set of unit Fubini rankings and their parking interpretation, they establish a complete characterization and enumeration for the total number of Boolean intervals in the weak Bruhat order of the symmetric group on $n$ letters, and the total number of Boolean intervals of rank $k$ in the same. Motivated by these results, we establish the following.
@@
      The specialization of $m=n$ is provided in Theorem 5.2, recovering the fact that unit interval parking functions are enumerated by the Fubini numbers [21].
2. Background
@@
2.1. Rational parking functions
In this scenario, there are $n$ cars and $m\geq n$ parking spots. A preference list $\alpha\in[m]^{n}$ permitting $n$ cars to park in $m$ spots utilizing the standard parking procedure is called a rational parking function of order $(n,m)$. Throughout, we let $\mathrm{PF}_{n,m}$ denote the set of rational parking functions of order $(n,m)$. As the order of the parking function is presented as a subscript, we henceforth omit stating the order of a rational parking function as it will be clear from context. Recall, that (1.1) gives $\left|\mathrm{PF}_{n,m}\right|=\left(m+1-n\right)(m+1)^{n-1}$. Moreover, if $n=m$, then $\mathrm{PF}_{n,n}=\mathrm{PF}_{n}$, and thus $\mathrm{PF}_{n}\subseteq\mathrm{PF}_{n,m}$, for all $m\geq n$. In this way, rational parking functions generalize parking functions.
2.2. Interval parking functions
We begin by remarking that in [12], interval parking functions were only defined for $m=n$. In this setting, interval parking functions are characterized as a pair $(\alpha,\beta)\in\mathrm{PF}_{n}\times[n]^{n}$, where $\alpha=(a_{1},a_{2},\ldots,a_{n})\in\mathrm{PF}_{n}$ and $\beta=(b_{1},b_{2},\ldots,b_{n})\in[n]^{n}$ satisfies $b_{i}\geq a_{i}$ for all $1\leq i\leq n$ [12, Proposition 3.2]. Given $\alpha\in\mathrm{PF}_{n}$ and $\beta\in[n]^{n}$, car $i$ enters the street and using the standard parking procedure, attempts to park between spots $a_{i}$ and $b_{i}$ (inclusive), which we refer to as its preferred interval. If all cars can park within their preferred interval utilizing the standard parking procedure, then the pair $(\alpha,\beta)$ is an interval parking function. We let $\mathrm{\mathrm{IPF}}_{n}$ denote the set of interval parking functions of length $n$ and recall (1.2) gives $|\mathrm{\mathrm{IPF}}_{n}|=n!(n+1)^{n-1}$.
@@ 2.2. Interval parking functions
2.3. Unit interval parking functions
Hadaway and Harris introduced unit interval parking functions, which are a subset of $\mathrm{PF}_{n}$ in which all cars park in their preferred spot or one spot further down the street [21]. Let $\mathrm{UPF}_{n}$ denote the set of unit interval parking functions of length $n$. Note, $(1,2,3,4,5)$, $(1,1,3,4,5)$, $(1,1,2,4,5)$ are elements of $\mathrm{UPF}_{5}$, whereas $(1,1,1,1,1)$ is not a unit interval parking function as car three cannot park in its preferred spot nor the one after it.
@@ 2.3. Unit interval parking functions
Theorem 2.2.
[21, Theorem 5.12] If $n\geq 1$, then
$|\mathrm{UPF}_{n}|$ $=\mathrm{Fub}_{n}$ (2.1)
@@
To establish Theorem 2.2, Hadaway and Harris proved that the set of unit interval parking functions is in bijection with the set of Fubini rankings ([21, Theorem 5.12]). As Fubini rankings play a role in our work, we define them next.
Definition 2.3.
A Fubini ranking of length $n$ is a tuple $r=(r_{1},r_{2},\ldots,r_{n})\in[n]^{n}$ that records a valid ranking of $n$ competitors with ties allowed (i.e., multiple competitors can be tied and have the same rank). If $k$ competitors are tied for rank $i$, the $k-1$ subsequent ranks $i+1,i+2,\ldots,i+k-1$ are disallowed. [^1] As noted in [15]: “One noteworthy instance of the tie condition of a Fubini ranking took place at the men’s high jump event at the Summer 2020 Olympics [14]. In this competition, Mutaz Essa Barshim of Qatar and Gianmarco Tamberi of Italy led the final round. Both athletes cleared 2.37 meters, but neither of them cleared 2.39 meters. Upon being presented the option of a “jump-off” to determine the sole winner, they agreed to instead share the gold medal. The next best rank was held by Maksim Nedasekau of Belarus, who obtained the bronze medal.”
For example, $(8,1,2,5,6,2,2,6)$ is Fubini ranking, in which competitor one ranks eighth, competitor two ranks first, competitors three, six, and seven tie at rank two, competitor four ranks fifth, and competitors five and eight rank sixth. Note $(1,1,1,2)$ is not a Fubini ranking as the triple tie for rank one would disallow ranks 2 and 3, making rank 4 the next available. We let $\text{FR}_{n}$ denote the set of Fubini rankings of length $n$. We remark that Cayley [8] showed that $|\text{FR}_{n}|=\mathrm{Fub}_{n}$; this together with the bijection of Hadway and Harris, establishes (2.1).
3. Interval and $\ell$-interval rational parking functions
@@
3.1. Interval rational parking functions.
@@
$\sigma_{i}=\begin{cases}j&\mbox{if car $j$ parks in spot $i$ under $\alpha$}\\ 0&\mbox{if no car parks in spot $i$ under $\alpha$}.\end{cases}$
@@
Figure 3. The parking outcome of the interval rational parking function $(4,3,5,5,9)\in\mathrm{PF}_{5,10}$.
@@
$I(\alpha)$ =$=\{i\in[m]:\sigma_{i}\neq 0\mbox{ where }\mathcal{O}_{m}(\alpha)=(\sigma_{1},\sigma_{2},\ldots,\sigma_{m})\}.$ (3.1)
@@
Example 3.3 (Continuing Example 3.2).
@@
Proof.
For any $\alpha\in\mathrm{PF}_{n,m}$, let $I(\alpha)$ be as in (3.1). As we noted above, if $\alpha\in\mathrm{PF}_{n,m}$, then $|I(\alpha)|=n$. Note that if $i\in I(\alpha)$, then $\sigma_{i}=j$ for a unique $j\in[n]$. Then $b_{j}$, the $j$th entry of the tolerance vector $\beta$, must satisfy $i\leq b_{j}\leq m$. That gives $m-i+1$ options for the possible value of $b_{j}$. As the tolerances of cars are independent, we can take the product of the possible values the tolerance can be. Thus, there are
@@
total tolerances for each $\alpha\in\mathrm{PF}_{n,m}$. The result follows from taking the sum of this product over all possible rational parking functions $\alpha\in\mathrm{PF}_{n,m}$. ∎
By setting $m=n$ in Theorem 3.4, we recover (1.2). We state and prove this below.
@@
[12, Proposition 3.2] If $n\geq 1$, then $|\mathrm{\mathrm{IPF}}_{n}|=n!(n+1)^{n-1}$.
Proof.
@@
$|\mathrm{\mathrm{IPF}}_{n}|=\sum_{\alpha\,\in\,\mathrm{PF}_{n}}\left(\prod_{i\,\in\,I(\alpha)}(n-i+1)\right)=\sum_{\alpha\,\in\,\mathrm{PF}_{n}}\left(\prod_{i\,\in\,[n]}(n-i+1)\right)=\sum_{\alpha\,\in\,\mathrm{PF}_{n}}n!=n!\,(n+1)^{n-1}.\qed$
3.2. $\ell-$interval parking functions

### v5

@@ Abstract
1 Introduction
Throughout, we let =$\mathbb{N}=\{0,1,2,3,\ldots\}$, +=$\mathbb{N}^{+}=\{1,2,3,\ldots\}$, and for $n\in\mathbb{N}^{+}$, we let $[n]=\{1,2,\ldots,n\}$. We begin by recalling a definition of parking functions, as introduced by Konheim and Weiss (1966). With $n\in\mathbb{N}^{+}$, a tuple $\alpha=(a_{1},a_{2},\ldots,a_{n})\in[n]^{n}$ is a parking function of length $n$ if its nondecreasing rearrangement $\alpha^{\prime}=(a_{1}^{\prime},a_{2}^{\prime},\ldots,a_{n}^{\prime})$ satisfies $a_{i}^{\prime}\leq i$ for all $i\in[n]$. For example, $\alpha=(3,1,1,3,4,5)$ is a parking function of length six as its nondecreasing rearrangement $\alpha^{\prime}=(1,1,3,3,4,5)$ satisfies the required inequality conditions. However, $\alpha=(2,3,3,3)$ is not a parking function, as its nondecreasing rearrangement $\alpha^{\prime}=(2,3,3,3)$ does not satisfy the inequality conditions: In particular, $a_{1}=2$ which is not less than or equal to $1$. Let $\mathrm{PF}_{n}\subseteq[n]^{n}$ denote the set of parking functions of length $n$. Several authors have provided proofs that $|\mathrm{PF}_{n}|=(n+1)^{n-1}$, see Konheim and Weiss (1966); Pyke (1959); Stanley (2016a).
@@ 1 Introduction
Figure 1: The parking outcome of the parking function $\alpha=(3,1,1,3,4,5)\in\mathrm{PF}_{6}$.
Parking functions have numerous applications and connections to many fields in mathematics, including computing volumes of flow polytopes, hyperplane arrangements, and the sorting algorithms Beck et al. (2015); Benedetti et al. (2019); Harris et al. (2024). Work on parking functions includes the study of discrete statistics Adeniran and Pudwell (2023); Gessel and Seo (2004/06); Schumacher (2018), families of parking functions with particular desirable properties Aguillon et al. (2023); Elder et al. (2023b), and many generalizations of parking functions Colaric et al. (2021); Colmenarejo et al. (2021); Franks et al. (2024); Harris et al. (2023). Our work is motivated by two generalizations of parking functions, which we refer to as “rational” parking functions and interval parking functions. We describe those next.
First we consider the situation in which there are $n\in\mathbb{N}^{+}$ cars and $m\geq n$ parking spots on a one-way street. In this case, a preference $\alpha$ is an element of $[m]^{n}$, and the set $\mathrm{PF}_{n,m}$ consists of the preferences allowing all $n$ cars to park under the standard parking procedure. This generalization of parking functions is known as $(m,n)$-parking functions. In the case of $n$ and $m$ being coprime, $(m,n)$-parking functions are referred to as rational parking functions, see Armstrong et al. (2016). In the current work, we simply refer to $(m,n)$-parking functions as rational parking functions without requiring $m$ and $n$ to be coprime. These parking functions were enumerated during the onset of parking function studies (Konheim and Weiss, 1966, Lemma 2) and, implicitly, in the statistical work of Pyke (1959), giving
$|\mathrm{PF}_{n,m}|=\left(m+1-n\right)(m+1)^{n-1}.$ (1)
Another generalization of parking functions are interval parking functions which describe the case where cars have an interval of the street in which they want to park. Formally, interval parking functions are defined by a pair $(\alpha,\beta)$, where $\alpha=(a_{1},a_{2},\ldots,a_{n})\in\mathrm{PF}_{n}$ and $\beta=(b_{1},b_{2},\ldots,b_{n})\in[n]^{n}$, satisfy, for all $i\in[n]$, the inequality $a_{i}\leq b_{i}\leq n$. For $i\in[n]$, we call $a_{i}$ the initial preference of car $i$, $b_{i}$ the tolerance of car $i$ (as it is the highest numbered parking spot in which car $i$ would tolerate parking), and the set $\{a_{i},a_{i}+1,\ldots,b_{i}\}$ the preference interval for car $i$. Throughout, we refer to $\beta$ as the tolerance vector. If, for each $i\in[n]$, car $i$ is able to park in its preference interval following the standard parking procedure, then the pair $(\alpha,\beta)$ is called an interval parking function. We let $\mathrm{\mathrm{IPF}}_{n}$ denote the set of interval parking functions of length $n$. Colaric, DeMuse, Martin, and Yin established (Colaric et al., 2021, Proposition 3.2):
$|\mathrm{\mathrm{IPF}}_{n}|=n!(n+1)^{n-1}.$ (2)
@@
In the present work, we begin by generalizing interval parking functions to their rational analogue, which we call interval rational parking functions. Here, there are $n$ cars and $m\geq n$ parking spots on the street. In this case, we consider the pair of tuples $(\alpha,\beta)$, where $\alpha=(a_{1},a_{2},\ldots,a_{n})\in\mathrm{PF}_{n,m}$ and $\beta=(b_{1},b_{2},\ldots,b_{n})\in[m]^{n}$ satisfying $a_{i}\leq b_{i}\leq m$. If, for each $i\in[n]$, car $i$ is able to park in one of the spots numbered $a_{i},a_{i}+1,\ldots,b_{i}$ (following the standard parking procedure), then $(\alpha,\beta)$ is an interval rational parking function. We let $\mathrm{\mathrm{IPF}}_{n,m}$ denote the set of interval rational parking functions with $n$ cars and $m$ spots. For example, $((3,5,7,1,1),(3,6,8,1,4))\in\mathrm{\mathrm{IPF}}_{5,8}$ is a rational interval parking function with five cars and eight parking spots whose parking outcome is illustrated in Figure 2. However, $((3,5,7,1,1),(3,6,8,1,1))$ is not an interval rational parking function for any $\mathrm{\mathrm{IPF}}_{5,m}$ with $m\geq 8$, as car $5$ would fail to park in the first spot, which is the only spot in its preference interval. In our first result, we establish a formula for $|\mathrm{\mathrm{IPF}}_{n,m}|$ as a sum over rational parking functions involving the parking outcome, i.e. the order in which the cars park on the street (Theorem 3.4). Specializing $m=n$ recovers the formula in Equation (2) for the number of interval parking functions.
@@
Figure 2: The parking outcome of the interval rational parking function $((3,5,7,1,1),(3,6,8,1,4))\in\mathrm{\mathrm{IPF}}_{5,8}$.
@@
Fubini rankings and unit interval parking functions have received much recent attention in the literature. In (Bradt et al., 2024, Theorem 2.9), the authors give a complete characterization of unit interval parking functions. In Chaves Meyles et al. (2023), the authors establish that unit interval parking functions with $k$ unlucky cars (cars that do not park in their preference) are in bijection with the $k$-dimensional faces of the permutohedron of order $n$. In Elder et al. (2023a), the authors consider the set of unit Fubini rankings, which is the intersection of the set of unit interval parking functions of length $n$ and the set of Fubini rankings. Through the set of unit Fubini rankings and their parking interpretation, they establish a complete characterization and enumeration for the total number of Boolean intervals in the weak Bruhat order of the symmetric group on $n$ letters, and the total number of Boolean intervals of rank $k$ in the same. Motivated by these results, we establish the following.
@@
      The specialization of $m=n$ is provided in Theorem 5.2, recovering the fact that unit interval parking functions are enumerated by the Fubini numbers Hadaway (2022).
2 Background
@@
2.1 Rational parking functions
In this scenario, there are $n$ cars and $m\geq n$ parking spots. A preference list $\alpha\in[m]^{n}$ permitting $n$ cars to park in $m$ spots utilizing the standard parking procedure is called a rational parking function of order $(n,m)$. Throughout, we let $\mathrm{PF}_{n,m}$ denote the set of rational parking functions of order $(n,m)$. As the order of the parking function is presented as a subscript, we henceforth omit stating the order of a rational parking function as it will be clear from context. Recall, that Equation (1) gives $\left|\mathrm{PF}_{n,m}\right|=\left(m+1-n\right)(m+1)^{n-1}$. Moreover, if $n=m$, then $\mathrm{PF}_{n,n}=\mathrm{PF}_{n}$, and thus $\mathrm{PF}_{n}\subseteq\mathrm{PF}_{n,m}$, for all $m\geq n$. In this way, rational parking functions generalize parking functions.
2.2 Interval parking functions
We begin by remarking that in Colaric et al. (2021), interval parking functions were only defined for $m=n$. In this setting, interval parking functions are characterized as a pair $(\alpha,\beta)\in\mathrm{PF}_{n}\times[n]^{n}$, where $\alpha=(a_{1},a_{2},\ldots,a_{n})\in\mathrm{PF}_{n}$ and $\beta=(b_{1},b_{2},\ldots,b_{n})\in[n]^{n}$ satisfies $b_{i}\geq a_{i}$ for all $1\leq i\leq n$ (Colaric et al., 2021, Proposition 3.2). Given $\alpha\in\mathrm{PF}_{n}$ and $\beta\in[n]^{n}$, car $i$ enters the street and using the standard parking procedure, attempts to park between spots $a_{i}$ and $b_{i}$ (inclusive), which we refer to as its preferred interval. If all cars can park within their preferred interval utilizing the standard parking procedure, then the pair $(\alpha,\beta)$ is an interval parking function. We let $\mathrm{\mathrm{IPF}}_{n}$ denote the set of interval parking functions of length $n$ and recall Equation (2) gives $|\mathrm{\mathrm{IPF}}_{n}|=n!(n+1)^{n-1}$.
@@ 2.2 Interval parking functions
2.3 Unit interval parking functions
Hadaway and Harris introduced unit interval parking functions, which are a subset of $\mathrm{PF}_{n}$ in which all cars park in their preferred spot or one spot further down the street Hadaway (2022). Let $\mathrm{UPF}_{n}$ denote the set of unit interval parking functions of length $n$. Note, $(1,2,3,4,5)$, $(1,1,3,4,5)$, $(1,1,2,4,5)$ are elements of $\mathrm{UPF}_{5}$, whereas $(1,1,1,1,1)$ is not a unit interval parking function as car three cannot park in its preferred spot nor the one after it.
@@ 2.3 Unit interval parking functions
Theorem 2.2
(Hadaway, 2022, Theorem 5.12) If $n\geq 1$, then
$|\mathrm{UPF}_{n}|=\mathrm{Fub}_{n}$ (3)
@@
To establish Theorem 2.2, Hadaway and Harris proved that the set of unit interval parking functions is in bijection with the set of Fubini rankings (Hadaway, 2022, Theorem 5.12). As Fubini rankings play a role in our work, we define them next.
Definition 2.3
A Fubini ranking of length $n$ is a tuple $r=(r_{1},r_{2},\ldots,r_{n})\in[n]^{n}$ that records a valid ranking of $n$ competitors with ties allowed (i.e., multiple competitors can be tied and have the same rank). If $k$ competitors are tied for rank $i$, the $k-1$ subsequent ranks $i+1,i+2,\ldots,i+k-1$ are disallowed. [^1] As noted in Elder et al. (2023a): “One noteworthy instance of the tie condition of a Fubini ranking took place at the men’s high jump event at the Summer 2020 Olympics Committee . In this competition, Mutaz Essa Barshim of Qatar and Gianmarco Tamberi of Italy led the final round. Both athletes cleared 2.37 meters, but neither of them cleared 2.39 meters. Upon being presented the option of a “jump-off” to determine the sole winner, they agreed to instead share the gold medal. The next best rank was held by Maksim Nedasekau of Belarus, who obtained the bronze medal.”
For example, $(8,1,2,5,6,2,2,6)$ is Fubini ranking, in which competitor one ranks eighth, competitor two ranks first, competitors three, six, and seven tie at rank two, competitor four ranks fifth, and competitors five and eight rank sixth. Note $(1,1,1,2)$ is not a Fubini ranking as the triple tie for rank one would disallow ranks 2 and 3, making rank 4 the next available. We let $\mathrm{FR}_{n}$ denote the set of Fubini rankings of length $n$. We remark that Cayley (2009) showed that $|\mathrm{FR}_{n}|=\mathrm{Fub}_{n}$; this together with the bijection of Hadway and Harris, establishes Equation (3).
3 Interval and $\ell$-interval rational parking functions
@@
3.1 Interval rational parking functions.
@@
$\sigma_{i}=\begin{cases}j\mbox{if car $j$ parks in spot $i$ under $\alpha$}\\ 0\mbox{if no car parks in spot $i$ under $\alpha$}.\end{cases}$
@@
Figure 3: The parking outcome of the interval rational parking function $(4,3,5,5,9)\in\mathrm{PF}_{5,10}$.
@@
$I(\alpha)=\{i\in[m]:\sigma_{i}\neq 0\mbox{ where }\mathcal{O}_{m}(\alpha)=(\sigma_{1},\sigma_{2},\ldots,\sigma_{m})\}.$ (5)
@@
Example 3.3 (Continuing Example 3.2)
@@
Proof 3.5.
For any $\alpha\in\mathrm{PF}_{n,m}$, let $I(\alpha)$ be as in Equation (5). As we noted above, if $\alpha\in\mathrm{PF}_{n,m}$, then $|I(\alpha)|=n$. Note that if $i\in I(\alpha)$, then $\sigma_{i}=j$ for a unique $j\in[n]$. Then $b_{j}$, the $j$th entry of the tolerance vector $\beta$, must satisfy $i\leq b_{j}\leq m$. That gives $m-i+1$ options for the possible value of $b_{j}$. As the tolerances of cars are independent, we can take the product of the possible values the tolerance can be. Thus, there are
@@
total tolerances for each $\alpha\in\mathrm{PF}_{n,m}$. The result follows from taking the sum of this product over all possible rational parking functions $\alpha\in\mathrm{PF}_{n,m}$.
By setting $m=n$ in Theorem 3.4, we recover Equation (2). We state and prove this below.
@@
(Colaric et al., 2021, Proposition 3.2) If $n\geq 1$, then $|\mathrm{\mathrm{IPF}}_{n}|=n!(n+1)^{n-1}$.
Proof 3.7.
@@
$|\mathrm{\mathrm{IPF}}_{n}|=\sum_{\alpha\,\in\,\mathrm{PF}_{n}}\left(\prod_{i\,\in\,I(\alpha)}(n-i+1)\right)=\sum_{\alpha\,\in\,\mathrm{PF}_{n}}\left(\prod_{i\,\in\,[n]}(n-i+1)\right)=\sum_{\alpha\,\in\,\mathrm{PF}_{n}}n!=n!\,(n+1)^{n-1}.$
3.2 $\ell-$interval parking functions